6. The recipe for cell culture media is: 10 % serum, 1x antibiotics and 2 mM Glutamine in basal media. You are given the foll please help solve this and show all work clearly, thanks
Answer #1

6. The calculations can be made using the dilution law

N1V1 = N2V2

N1 = Initial concentration of the component

V1 = Initial volume of the solution

N2 = Final concentration of the component

V2 = Final volume of the solution

Given that the recipe for the cell culture is

10 % serum

1x antibiotics

2 mM Glutamine

These are the final concentrations of serum, antibodies and glutamine in the final 500 mL cell culture media.

Given stock solution are

serum - 100 %

antibiotics - 100x

glutamine - 200 mM

Now we should calculate the initial volumes of stocks to prepare 500mL cell culture media

1. Serum

N1V1 = N2V2

100 % x V1 = 10 % x 500 mL

V1 = 5000 / 100 = 50 mL of serum

2. antibiotics

N1V1 = N2V2

(100x) x V1 = (1x) x 500 mL

V1 = 500/100 = 5 mL of antibiotics

3. Glutamine

N1V1 = N2V2

200 mM x V1 = 2 mM x 500 mL

V1 = 1000 / 200 = 5 mL

4. Basal media

These three are taken in basal media. Hence the cell culture media should be make up to 500 mL with basal media.

Thus the volume of basal media = Volume of final solution - Volume of other components

The volume of basal media = 500 - (50 + 5 + 5) = 500 - 60 = 440 mL

Thus we can prepare 500 mL of cell culture media by taking 440 mL of basal media, 50 mL of stock serum, 5 mL of stock antibiotics and 5 mL of stock glutamine